# Inverse-Squeezing Kennedy Receiver for Near-Helstrom Discrimination of Displaced-Squeezed BPSK

### v1

@@ III Inverse-squeezing Kennedy Receiver under Ideal Conditions
$\hat{i}(n)=\begin{cases}1,&P(i=1|n)\geq P(i=0|n),\\ 0,&\text{otherwise}.\end{cases}$ (16)
Assuming equal priors ($\pi_{0}=\pi_{1}=1/2$), the log-likelihood ratio is:
$L(n)=\ln\frac{P(n|i=0)}{P(n|i=1)}=\ln\frac{\text{Poiss}\left(n;\mu_{0}\right)}{\text{Poiss}\left(n;\mu_{1}\right)}=-\mu_{0}+\mu_{1}+n\ln\left(\frac{\mu_{0}}{\mu_{1}}\right).$ (17)
Since $L(n)$ is a monotonic function of $n$, the MAP decision (Eq. (16)) is equivalent to a threshold detection:
$\hat{i}(n)=\begin{cases}1,&n\geq n_{\text{th}}^{*},\\ 0,&\text{otherwise}.\end{cases}$ (18)
The optimal threshold $n_{\text{th}}^{*}$ corresponds to the zero-crossing of the log-likelihood ratio:
$n_{\text{th}}^{*}=\left\lceil\frac{\mu_{1}-\mu_{0}}{\ln\mu_{1}-\ln\mu_{0}}\right\rceil.$ (19)
Under ideal conditions where $\mu_{0}=0$, we have $\lim_{\mu_{0}\to 0}\frac{\mu_{1}-\mu_{0}}{\ln\mu_{1}-\ln\mu_{0}}\to 0^{+}$; thus, $n_{\text{th}}^{*}=1$. This corresponds to a standard on-off detection strategy, implying that a simple single-photon detector (SPD) suffices to achieve PNR performance in this ideal limit. It is important to emphasize, however, that the introduction of the PNRD is not intended to improve the limit of “ideal nulling with matched squeezing.” Rather, it addresses practical system imperfections: once squeezing mismatch or detector imperfections (efficiency/dark counts) occur, $\zeta_{0}$ deviates from the vacuum state, yielding a non-zero $P(n\geq 1|i=0)$. In such cases, the optimal decision is no longer a simple on-off rule, and the PNRD provides the necessary statistical information to implement a generalized MAP or optimal threshold strategy. Quantitative analysis of these non-ideal factors is presented in subsequent sections.
The false alarm probability $P_{\text{FA}}$ (deciding 1 when 0 is sent) and the miss detection probability $P_{\text{Mi}}$ (deciding 0 when 1 is sent) are:
$P_{\text{FA}}=P\left(n\geq n_{\text{th}}^{*}\mid i=0\right),\quad P_{\text{Mi}}=P\left(n<n_{\text{th}}^{*}\mid i=1\right).$ (20)
Therefore, under equal priors and ideal conditions ($n_{\text{th}}^{*}=1$), the average error probability of the IS-Kennedy is:
$P_{\text{err}}^{\text{IS-K},\text{ideal}}$ $=\frac{1}{2}\left(P_{\text{FA}}+P_{\text{Mi}}\right)$ (21)
$=\frac{1}{2}\left[1-P(n=0|i=0)+P(n=0|i=1)\right]$
$=\frac{1}{2}\exp\left(-|2\gamma|^{2}\right)=\frac{1}{2}\exp\left(-4N_{\text{eff}}\right).$
Substituting the optimal energy allocation condition $N_{\text{eff}}=N(N+1)$, we obtain:
$P_{\text{err}}^{\text{IS-K},\text{ideal}}=\frac{1}{2}\exp\left[-4N(N+1)\right].$ (22)
Notably, we observe that:
$P_{\text{err}}^{\text{IS-K},\text{ideal}}(N)=P_{\text{err}}^{\text{K},\text{ideal}}(N_{\text{eff}}),$ (23)
where $P_{\text{err}}^{\text{K},\text{ideal}}$ is the error probability of the conventional Kennedy receiver under ideal conditions [15]. This implies that, under matched squeezing, the IS-Kennedy is equivalent to mapping an S-BPSK signal with mean photon number $N$ to a C-BPSK signal with mean photon number $N_{\text{eff}}$ via the IS module, followed by Kennedy nulling and on-off detection.
For S-BPSK, the Helstrom bound in Eq. (9) and the IS-Kennedy error probability in Eq. (22) satisfy the relation:
$\frac{P_{\text{err}}^{\text{IS-K},\text{ideal}}}{P_{\text{HB}}^{\text{DSS}}}=\frac{\exp\left\{-4N\left(N+1\right)\right\}}{1-\sqrt{1-\exp\left\{-4N(N+1)\right\}}}.$ (24)
In the low-energy limit $N\to 0$, both $P_{\text{err}}^{\text{IS-K}}$ and $P_{\text{HB}}^{\text{DSS}}$ approach $0.5$, and their ratio approaches $1$, corresponding to random guessing. In the high-energy regime $N\gg 1$, the ratio approaches $2$ (i.e., $3$ dB), yielding:
$P_{\text{err}}^{\text{IS-K},\text{ideal}}\simeq 2P_{\text{HB}}^{\text{DSS}}.$ (25)
Although the IS-Kennedy remains suboptimal with respect to the S-BPSK Helstrom bound, satisfying $P_{\text{err}}^{\text{IS-K}}\in\left[P_{\text{HB}}^{\text{DSS}},2P_{\text{HB}}^{\text{DSS}}\right]$ for all $N$, it provides a substantial advantage over C-BPSK channels. Specifically, for the same mean photon number $N$, the proposed IS-Kennedy achieves an error probability strictly lower than the Helstrom bound $P_{\text{HB}}^{\text{CS}}$ of a C-BPSK channel. For instance, around $N\approx 0.4$ in Fig. 4, the IS-Kennedy performance lies between the DSS Helstrom bound and the coherent-state Helstrom bound.
Figure 4: Performance of the IS-Kennedy under ideal conditions. (a) Error probability of the IS-Kennedy. (b) Ratio (in dB) of the benchmarks $P_{\text{HB}}^{\text{CS}}$, $P_{\text{SQL}}^{\text{CS}}$, $P_{\text{HB}}^{\text{DSS}}$, $P_{\text{SQL}}^{\text{DSS}}$, and $P_{\text{err}}^{\text{K},\text{ideal}}$ relative to $P_{\text{err}}^{\text{IS-K},\text{ideal}}$.
Figure 4(a) shows the IS-Kennedy error probability as a function of the mean photon number $N$ under ideal conditions, alongside the Kennedy receiver and benchmarks for C-BPSK and S-BPSK discrimination. The IS-Kennedy surpasses the C-BPSK SQL, the S-BPSK SQL, and the C-BPSK Helstrom bound at approximately $N\approx 0.21$, $0.30$, and $0.40$, respectively. The performance advantage over these benchmarks increases with $N$, consistent with the fact that $P_{\text{err}}^{\text{IS-K}}=\frac{1}{2}\exp[-4N(N+1)]$ features an approximately quadratic exponent $N(N+1)$, leading to a steeper decay than the linear scaling of coherent-state benchmarks. The dB ratios summarized in Fig. 4(b) quantify these advantages. At $N=1.0$, the IS-Kennedy achieves gains of $21.3$ dB, $11.4$ dB, and $14.4$ dB relative to the C-BPSK SQL, the S-BPSK SQL, and the C-BPSK Helstrom bound, respectively. Additionally, the IS-Kennedy significantly outperforms the conventional Kennedy receiver, exhibiting a $17.4$ dB gain at $N=1.0$. Meanwhile, it remains $3$ dB above the S-BPSK Helstrom bound, staying within a constant factor of the ultimate quantum limit. These results demonstrate that the IS-Kennedy delivers robust performance over a broad photon-number regime under ideal conditions.
IV IS-Kennedy Receiver with an Imperfect PNR Detector
The previous section presented the closed-form error probability of the IS-Kennedy receiver under ideal detection conditions, pointing out that with perfect nulling and matched squeezing, the optimal decision rule degenerates to a simple on–off strategy. However, in practical implementations, PNR detectors possess finite resolution (maximum resolvable photon number, M) and suffer from non-ideal factors such as limited detection efficiency and dark counts (background counts). In this section, while maintaining the assumption of an ideal IS-Kennedy processing module (i.e., retaining $|\zeta_{0}=\ket{0}\bra{0}$ and $|\zeta_{1}=\ket{2\gamma}\bra{2\gamma}$), we analyze the impact of imperfect PNR detection on the error probability and derive the expression for the optimal threshold (MAP-equivalent) decision and the corresponding error probability.
Considering an imperfect detection efficiency $\eta\in(0,1]$ and a mean dark count rate $\nu\geq 0$, the POVM operators are described by [2, 12]:
$|.\Pi_{n}^{\left(\eta,\nu\right)}=\mathrm{e}^{-\nu}\sum_{l=0}^{n}\sum_{k=n-l}^{\infty}\frac{\nu^{l}}{l!}\binom{k}{n-l}\eta^{n-l}(1-\eta)^{k-(n-l)}\ket{k}\bra{k}$. (26)
Consequently, the probability of detecting $n$ photons (Eq. (14)) is reformulated as:
$P(n|i)=\text{Tr}\left\{\zeta_{i}\Pi_{n}^{\left(\eta,\nu\right)}\right\}=\left\{\begin{aligned} &\text{Poiss}\left(n;\eta\mu_{0}+\nu\right),&i=0,\\ &\text{Poiss}\left(n;\eta\mu_{1}+\nu\right),&i=1,\end{aligned}\right$. (27)
where $\mu_{0}=0$ and $\mu_{1}=\left|2\gamma\right|^{2}=4N_{\text{eff}}$. The log-likelihood ratio (Eq. (17)) is updated to:
$L(n)=-\eta\mu_{0}+\eta\mu_{1}+n\ln\frac{\eta\mu_{0}+\nu}{\eta\mu_{1}+\nu},$ (28)
which remains a monotonic function of $n$. Therefore, the MAP criterion is equivalent to a threshold detection scheme. Taking the PNR detector’s resolution M into account (denoted as PNR(M)), the optimal threshold (Eq. (19)) involves a truncation at M:
$n_{\text{th}}^{*}=\min\left\{\left\lceil\frac{\eta\mu_{1}-\eta\mu_{0}}{\ln\left(\eta\mu_{1}+\nu\right)-\ln\left(\eta\mu_{0}+\nu\right)}\right\rceil,\text{M}\right\}=\min\left\{\left\lceil\frac{4\eta N_{\text{eff}}}{\ln\left(4\eta N_{\text{eff}}+\nu\right)-\ln\nu}\right\rceil,\text{M}\right\}.$ (29)
From the above equation, the adaptive behavior of the IS-Kennedy receiver against different non-ideal factors can be observed:
  • Presence of dark counts ($\nu>0$): The denominator decreases, resulting in an optimal threshold $n_{\text{th}}^{*}>1$. This implies that the PNR detector can filter out background noise by raising the decision threshold, demonstrating its robustness advantage over a standard single-photon detector (SPD).
  • Negligible dark counts ($\nu\to 0$): The denominator approaches infinity, leading to $n_{\text{th}}^{*}\to 1$. In this limit, the optimal strategy reverts to the standard on–off decision. This indicates that in a pure loss channel ($\eta<1,\nu=0$), the PNR detector offers no performance gain over an SPD, as simple presence/absence detection is optimal.
Figure 5: Performance analysis of the IS-Kennedy receiver under imperfect detection conditions. (a) Error probability of IS-Kennedy with SPD for varying detection efficiencies $\eta\in\left\{1.0,0.8,0.5\right\}$ and extremely small dark counts ($\nu=10^{-9}$). (b) Error probability of IS-Kennedy with PNR(2) under different dark count rates $\nu$ (with $\eta=1.0$). The “step-like” behavior in the error curves arises from the discrete nature of the optimal decision threshold. (c) The optimal detection threshold $n_{\text{th}}^{*}$ as a function of signal energy for different dark count rates, calculated via Eq. (29). The jumps in threshold correspond exactly to the inflection points observed in panels (b) and (d). (d) The performance ratio (in dB) of the IS-Kennedy with PNR(10) relative to the S-BPSK SQL. The oscillatory behavior reflects the discrete adjustment of $n_{\text{th}}^{*}$ to balance signal detection against dark count noise.
Figure 5 illustrates the performance of the IS-Kennedy receiver with imperfect detectors. As shown in Fig. 5(b) and (d), the error probability and the gain relative to the SQL exhibit a characteristic oscillatory or “step-like” behavior when dark counts are significant. This phenomenon is intrinsic to the discreteness of PNR detection. As the signal energy $N$ increases, the optimal likelihood ratio varies continuously, whereas the physical decision threshold $n_{\text{th}}^{*}$ must remain an integer. Consequently, $n_{\text{th}}^{*}$ stays constant over intervals of $N$ (see Fig. 5(c)) before discretely jumping to the next integer level. These jumps allow the receiver to adaptively filter out higher dark-count rates, maintaining robust performance even when $\nu$ is large (e.g., $\nu=10^{-2}$).
Notably, Fig. 5(a) reveals that while the detection efficiency $\eta$ impacts the rate of error decay (the slope), the saturation error floor at high energies is dictated solely by the dark count rate $\nu$. In the high-energy regime ($N\gg 1$), the miss detection probability (Eq. (20)) rapidly approaches zero. Consequently, the error probability (saturation floor) is dominated by the false alarm probability, with the optimal threshold saturating at $n_{\text{th}}^{*}=\text{M}$:
$P_{\text{sat}}\approx\frac{1}{2}P_{\text{FA}}=\frac{1}{2}\mathrm{e}^{-\nu}\sum_{n=\text{M}}^{+\infty}\frac{\nu^{n}}{n!}\xrightarrow{\text{Taylor series}}\frac{1}{2}\left(1-\nu+\cdots\right)\sum_{n=\text{M}}^{+\infty}\frac{\nu^{n}}{n!}.$ (30)
Given that the dark count rate of practical PNR detectors is typically much less than unity ($\nu\ll 1$[6], the dominant contribution to the summation comes from the first term, $n=\text{M}$. Thus, the saturation error probability can be approximated as:
$P_{\text{sat}}\approx\frac{1}{2}\frac{\nu^{\text{M}}}{\text{M}!}.$ (31)
This expression indicates that the saturation error probability of the IS-Kennedy receiver is independent of detection efficiency $\eta$ and mean photon number $N$, depending instead on the PNR detector’s resolution M and dark count rate $\nu$. Therefore, the existence of any non-zero dark count, however small, inevitably leads to an error probability saturation. Furthermore, in conjunction with Eq. (23), since the advantage of the IS-Kennedy equipped with an SPD lies in converting squeezing resources into effective energy gain rather than suppressing noise floor, its saturation error probability is identical to that of the conventional Kennedy receiver (see Fig. 5(a)).
V IS-Kennedy Receiver with Squeezing Mismatch
In the preceding sections, we assumed that the IS module at the receiver could perfectly execute the S(r)S(-r) operation, thereby precisely converting the squeezing resources into an enhanced energy gain. However, in practical experimental setups, deviations between the receiver’s squeezing parameters and the input signal are inevitable due to factors such as optical loss, pump fluctuation, and phase locking errors. This phenomenon is referred to as squeezing mismatch (SM).
@@ V IS-Kennedy Receiver with Squeezing Mismatch
Without loss of generality, we adopt a simplified model where the transmitter performs perfect squeezing and the optical path is lossless, attributing all squeezing adaptation errors solely to the IS module. Let the squeezing parameter applied by the IS module be denoted as:
@@
[…]
Here, the composite squeezing operator is defined as $S_{\text{tot}}\triangleq S\left(z_{s}\right)S(r)$, which reduces to the identity operator $S_{\text{tot}}=\mathbb{I}$ in the ideal case. Thus, the essence of the mismatch is encapsulated within $S_{\text{tot}}$. Using the single-mode Gaussian operator decomposition, this operator can always be expressed as:
@@
We now calculate the photon number probability distribution of $\zeta_{i}$. Ignoring the rotation $R(\vartheta)$ which does not affect photon statistics, the equivalent output states derived from Eq. (34) are:
@@
[…]
@@
Consequently, the Poissonian distribution valid for the ideal IS-Kennedy (Eq. (14)) no longer applies. For an ideal PNR(M) detector (considering only finite resolution), the probability of detecting $n$ photons given symbol $i$ is:
$P(n|i)=\text{Tr}\left(\Pi_{n}\,\zeta_{i}\right)=\left\{\begin{aligned} &p_{n}^{\rm sv}(r_{m}),&i=0,\\ &p_{n}^{\rm dss}(2\gamma_{m},r_{m},\theta_{m}),&i=1,\end{aligned}\right$. (44)
where the corresponding distribution is used for $n\leq\text{M}-1$, and the truncation bin is defined as $P(n=\text{M}|i)=1-\sum_{n=0}^{\text{M}-1}P(n|i)$. Here, $p_{n}^{\rm sv}(r_{m})$ denotes the photon number distribution of the squeezed vacuum state (Eq. (56)), and $p_{n}^{\rm dss}(\gamma_{m},r_{m},\theta_{m})$ denotes that of the displaced squeezed vacuum state (Eq. (58)).
The introduction of mismatch results in $\zeta_{0}$ becoming a squeezed vacuum state (containing only even photon numbers), while $\zeta_{1}$ becomes a displaced squeezed vacuum state whose Fock distribution generally exhibits even-odd oscillations (see Fig. 6(d)). This implies that the likelihood ratio does not necessarily vary monotonically with the photon number $n$. Consequently, the MAP decision generally becomes a set-based decision (potentially involving a multi-threshold structure) rather than a simple single-threshold test. Therefore, the minimum error probability (Eq. (21)) is reformulated as:
$P_{\text{err}}^{\text{IS-K},\text{SM}}=1-\frac{1}{2}\sum_{n=0}^{\text{M}}\max\left\{P(n|i=0),\ P(n|i=1)\right\}.$ (45)
When employing an SPD, the observation yields only two outcomes: $n=0$ (no-click) and $n\geq 1$ (click). Under equal priors, the MAP decision is equivalent to a binary mapping of these outcomes. Since $P(0|0)\geq P(0|1)$ (as the exponent in the latter is $\leq 0$), the MAP rule maps to on–off detection: click decides symbol 1, no-click decides symbol 0. The false alarm probability is:
@@
These equations clearly reveal that the residual squeezing $r_{m}$ reduces the vacuum component of $\zeta_{0}$, thereby increasing the false alarm probability, while $\theta_{m}$ modifies the vacuum overlap of $\zeta_{1}$ through the exponential term.
Figure 6: Performance analysis of the IS-Kennedy receiver under inverse-squeezing mismatch conditions. (a) Error probability of IS-Kennedy with SPD. (b) Error probability of IS-Kennedy with PNR(3). (c) The performance ratio (in dB) of IS-Kennedy with PNR(10) relative to the S-BPSK SQL. (d) The Fock-basis population of $\zeta_{0},\ \zeta_{1}$ with inverse-squeezing mismatch of $\Delta r=0.02,\ \Delta\theta=0.03\pi$.
Furthermore, we consider the saturation characteristics in the case of amplitude-only mismatch ($\Delta r\neq 0,\Delta\theta=0$). As the signal energy tends to infinity ($N\to\infty$), the mean photon number of the signal state $\zeta_{1}$ far exceeds the resolution M of the PNR detector, leading to $P_{\text{Mi}}\to 0$. In this regime, the error probability of the IS-Kennedy is dominated by the false alarm probability. The MAP strategy is thus equivalent to a threshold decision with $n_{\text{th}}^{*}=\text{M}$: decide symbol 1 if the detected photon number exceeds M, and symbol 0 otherwise. Since $\zeta_{0}$ is a squeezed vacuum state with a non-zero photon number distribution only for even terms (see Eq. (56)), the saturation error probability is:

### v2

@@ II.1 Receiver under ideal condition
$\hat{i}(n)=\left\{\begin{aligned} &1,\quad P(i=1|n)\geq P(i=0|n),\\ &0,\quad\text{otherwise}.\end{aligned}\right$. (10)
where $\hat{i}$ is the decision of IS-Kennedy receiver. When the likelihood function $L(n)=\ln\frac{P(n|i=0)}{P(n|i=1)}$ is monotonic, the MAP decision is equivalent to a single-threshold decision[1], with the optimal threshold denoted as $n_{\text{th}}^{*}$. In addition, under ideal conditions, the optimal threshold is 1, which means that PNR detector does not provide a performance advantage over SPD in this case [1]. So the error probability of IS-Kennedy is given by:
$P_{\text{err}}^{\text{ISK, ideal}}$ $=\frac{1}{2}\left(P_{\text{FA}}+P_{\text{Mi}}\right)=\frac{1}{2}\left[P(n\geq n_{\text{th}}^{*}|i=0)+P(n<n_{\text{th}}^{*}|i=1)\right]$ (11)
$=P_{\text{err}}^{\text{K,ideal}}(N_{\text{eff}})=\frac{1}{2}\exp\left[-4N(N+1)\right],$
where $P_{\text{FA}}$ and $P_{\text{Mi}}$ are the false alarm probability (deciding 1 when 0 is sent) and the miss probability (deciding 0 when 1 is sent), respectively, and $P_{\text{err}}^{\text{K,ideal}}$ is the ideal probability of the Kennedy receiver [5]. From this, it can be seen that the IS-Kennedy converts the squeezing resources of S-BPSK signals into higher coherent energy advantage ($N\to N_{\text{eff}}=N(N+1)$) through the IS module, thereby achieving a lower error probability. Compared to quantum receiver for coherent state discrimination, the energy gain brought by the IS-Kennedy is:
$G=10\cdot\log_{10}\left(\frac{N_{\text{eff}}}{N}\right)=10\cdot\log_{10}\left(N+1\right)\,(\text{dB})$ (12)
II.2 Receiver under squeezing mismatch condition
@@
In practical experimental setups, deviations between the receiver’s squeezing parameters and the input signal are inevitable due to factors such as optical loss, pump fluctuation, and phase locking errors. This phenomenon is referred to as squeezing mismatch (SM). Without loss of generality, we adopt a simplified model where the transmitter performs perfect squeezing and the optical path is lossless, attributing all squeezing adaptation errors solely to the IS module. Let the squeezing parameter applied by the IS module be denoted as:
@@
Under mismatch conditions, the overall transformation of the receiver is given by:
@@
Here, the composite squeezing operator is defined as $S_{\text{tot}}\triangleq S\left(z_{s}\right)S(r)$, which reduces to the identity operator $S_{\text{tot}}=\mathbb{I}$ in the ideal case. Thus, the essence of the mismatch is encapsulated within $S_{\text{tot}}$.
Using the single-mode Gaussian operator decomposition, this operator can always be expressed as:
@@
We now calculate the photon number probability distribution of $\zeta_{i}^{\text{sm}}$. Ignoring the rotation $R(\vartheta)$ which does not affect photon statistics, the equivalent output states derived from Eq. (15) are:
@@
To utilize standard photon number distribution formulas, we rewrite $\zeta_{1}^{\text{sm}}$ as a displaced squeezed vacuum state:
@@
Consequently, the Poissonian distribution valid for the ideal IS-Kennedy (Eq. (8)) no longer applies. For an ideal PNR(M) detector (considering only finite resolution), the probability of detecting $n$ photons given hypothes $i$ is:
$P(n|i)=\text{Tr}\left(\Pi_{n}\,\zeta_{i}^{\text{sm}}\right)=\left\{\begin{aligned} &p_{n}^{\rm sv}(r_{m}),&i=0,\\ &p_{n}^{\rm dss}(2\gamma_{m},r_{m},\theta_{m}),&i=1,\end{aligned}\right$. (25)
where the corresponding distribution is used for $n\leq\text{M}-1$, and the truncation/saturation bin is defined as $P(n=\text{M}|i)=1-\sum_{n=0}^{\text{M}-1}P(n|i)$. Here, $p_{n}^{\rm sv}(r_{m})$ denotes the photon number distribution of the squeezed vacuum state (Eq. (32)), and $p_{n}^{\rm dss}(2\gamma_{m},r_{m},\theta_{m})$ denotes that of the displaced squeezed vacuum state (Eq. (34)).
The introduction of mismatch results in $\zeta_{0}^{\text{sm}}$ becoming a squeezed vacuum state (containing only even photon numbers), while $\zeta_{1}^{\text{sm}}$ becomes a displaced squeezed vacuum state whose Fock distribution generally exhibits even-odd oscillations (see Fig. 2(d)). This implies that the likelihood function $L(n)$ does not necessarily vary monotonically with the photon number $n$. Consequently, the MAP decision generally becomes a set-based decision (potentially involving a multi-threshold structure) rather than a simple single-threshold test. Therefore, the minimum error probability (Eq. (11)) is reformulated as:
$P_{\text{err}}^{\text{ISK},\text{sm}}=1-\frac{1}{2}\sum_{n=0}^{\text{M}}\max\left\{P(n|i=0),\ P(n|i=1)\right\}.$ (26)
When employing an SPD, the observation yields only two outcomes: $n=0$ (no-click) and $n\geq 1$ (click). Under equal priors, the MAP decision is equivalent to a binary mapping of these outcomes. Since $P(0|0)\geq P(0|1)$ (as the exponent in the latter is $\leq 0$), the MAP rule maps to on–off detection: click decides hypothes 1, no-click decides hypothes 0. The false alarm probability is:
@@
These equations clearly reveal that the residual squeezing $r_{m}$ reduces the vacuum component of $\zeta_{0}^{\text{sm}}$, thereby increasing the false alarm probability, while $\theta_{m}$ modifies the vacuum overlap of $\zeta_{1}^{\text{sm}}$ through the exponential term.
Furthermore, we consider the saturation characteristics in the case of amplitude-only mismatch ($\Delta r\neq 0,\Delta\theta=0$). As the signal energy tends to infinity ($N\to\infty$), the mean photon number of the signal state $\zeta_{1}^{\text{sm}}$ far exceeds the resolution M of the PNR detector, leading to $P_{\text{Mi}}\to 0$. In this regime, the error probability of the IS-Kennedy is dominated by the false alarm probability. The MAP strategy is thus equivalent to a threshold decision with $n_{\text{th}}^{*}=\text{M}$: decide hypothes 1 if the outcome falls into the saturation bin $n=\text{M}$, and hypothes 0 otherwise. Since $\zeta_{0}^{\text{sm}}$ is a squeezed vacuum state with a non-zero photon number distribution only for even terms (see Eq. (32)), the saturation error probability is: